Evaluate the integral of z tan z² dz


1
Expert's answer
2021-03-07T09:47:30-0500

"I= \int z(tanz^2) dz"


Let "z^2=t\Rightarrow 2zdz=dt\\"

"\Rightarrow zdz=\dfrac{dt}{2}"

Now, "I= \int tan(t)\times \dfrac{dt}{2}"

"\Rightarrow I=\dfrac{1}{2}\int tan(t)dt"

"\Rightarrow I= \dfrac{1}{2}\times ln|sec(t)|+C"


"\Rightarrow I= \dfrac{1}{2}" "ln |sec(z^2)|+C"


"\int ztan(z^2)dz=\dfrac{ln|sec(z^2)|}{2}+C"